# SonicMaster: Towards Controllable All-in-One Music Restoration and Mastering

### v2

@@ 5.1 Objective Evaluation
SonicMasterLarge (6/18) 0.095 0.754 0.380 7.627 0.121 1.251 0.368 7.477 0.101 1.002 0.374 7.552
Table 3: Objective scores: FAD, KL, SSIM, and PQ. For readability, KL are multiplied by 1000.
@@
Perceptual Quality Assessment: Table 5.1 reveals SonicMaster outperforms the degraded inputs in both PQ and KL. FAD is marginally higher than that of the degraded audio, yet markedly lower than the reconstructed baseline. Furthermore, SonicMaster achieves a significant increase in PQ, almost reaching the level of ground truth mastered reference. In SSIM, SonicMaster exhibits lower scores than degraded inputs but achieves superior performance compared to the reconstruction baseline.
Architecture Scaling Analysis: Tables 4.1, 5.1, 5.1, reveal interesting scaling dynamics. SonicMasterSmall performs comparably with SonicMasterLarge in all metrics, but slightly worse in Reverb, Clip, and Stereo. SonicMasterMedium (4MM-DiT/12DiT) performs slightly better than the SonicMasterSmall, but still lacks behind SonicMasterLarge in Clip. SonicMasterMedium (6MM-DiT/6DiT) performs the worst out of all variants across all metrics.
5.2 Ablation Studies
Audio Conditioning Duration: We evaluated SonicMasterLarge with different conditioning lengths (5s, 10s, 15s), finding comparable performance across configurations. The 10-second setting balances computational efficiency with temporal overlap for long-form processing.
Conditioning Strategy Analysis: The no-conditioning variant achieves optimal Boom correction (0.0658 absolute error) but poor Clip restoration (2.055 vs. standard variants), highlighting multi-modal guidance importance for challenging tasks.
Text Prompt Dependency:Inference without text prompts maintains comparable FAD, SSIM, and PQ but shows degraded KL divergence (0.917 vs. 0.696). Critical drops occur in Clip restoration (2.812 vs. 1.506) and Stereo processing (0.1416 vs. 0.1058), with elevated EQ errors. This confirms text conditioning enables targeted restoration rather than generic improvements.
@@ 5.2 Ablation Studies
Long-Form Audio Evaluation: Full-song evaluations confirm SonicMaster’s effectiveness, with substantial improvements in EQ-related metrics (Table 4.1) and most degradation functions (Table 5.1). Reverberation metrics show mixed results, likely due to increased complexity of spatial processing in extended musical contexts where room acoustics interact with diverse instrumental timbres and dynamic variations. SSIM and FAD decrease compared to degraded inputs, except for FAD in multi-degradation samples, indicating SonicMaster’s ability to handle compound degradations.
@@ 5.3 Piano Recordings Evaluation
To test SonicMaster generalization, we evaluate historical solo piano pieces [^3] http://research.spa.aalto.fi/publications/papers/dafx-babe2/ using established baselines: LTAS-EQ, BEHM-GAN (Moliner & Välimäki, 2023) model for bandwidth extension, and BABE/BABE-2 diffusion-based generative equalizers (Moliner et al., 2024; Moliner & Välimäki, 2023). BABE-2 represents a state-of-the-art specialized method for old recordings, it uses a diffusion prior to restore lost high frequencies and remove coloration, and has shown impressive improvements in archival music (Moliner et al., 2024). Despite lacking domain-specific training, SonicMaster came surprisingly close to these specialized baselines. In objective evaluations, SonicMaster restored samples achieved a PQ of $6.93$, nearly matching the $7.05$ obtained by BABE-2.
Category Text relevance Quality1 Quality2 Consistency Preference
EQ $5.11_{\pm 0.55}$ $4.55_{\pm 0.44}$ $1_{\pm 0.52} 5.05±0$$5_{\pm 0.42}$ $4.84_{\pm 0.49}$
Reverb $5.59_{\pm 0.57}$ $4.25_{\pm 0.68}$ $5.28_{\pm 0.57}$ $5.31_{\pm 0.54}$ $5.27_{\pm 0.75}$
Dynamics $4.56_{\pm 0.80}$ $4.09_{\pm 0.54}$ $5.00_{\pm 0.81}$ $5.00_{\pm 0.84}$ $4.91_{\pm 1.01}$
Amplitude $6.19_{\pm 0.66}$ $3.66_{\pm 0.43}$ $38_{\pm 0.77} 5.5$$59_{\pm 0.81}$ $5.75_{\pm 0.84}$
Stereo $5.50_{\pm 1.05}$ $4.62_{\pm 0.83}$ $5.50_{\pm 0.63}$ $5.38_{\pm 0.70}$ $5.25_{\pm 0.77}$
Mixed degradations $5.28_{\pm 0.51}$ $3.85_{\pm 0.88}$ $4.60_{\pm 0.59}$ $4.65_{\pm 0.62}$ $4.93_{\pm 0.57}$
Table 5: Listening study - SonicMaster’s performance on specific degradations – MOS 95% CI.
@@ 5.4 Subjective Evaluation
Table 5 shows results of the first subjective evaluation test. Text relevance ratings are highest in the Amplitude (6.19) and Reverb categories (5.59), indicating effective declipping, volume increase, and dereverberation. These two categories also show the highest improvement in quality, as well as the highest consistency and preference ratings. EQ shows the fourth-best text relevance, but the worst preference ratings. Overall, SonicMaster samples are rated higher in quality compared to inputs and preferred across the board. A paired $t$-test on Quality1 and Quality2 ratings shows statistically significant differences in Reverb, Amplitude, Stereo, and mixed degradations.
Category SonicMaster Mel2Mel + Diffwave (Kandpal et al., 2022) Text2FX (Chu et al., 2025)
Reverberation 128/130 (98.5%) 2/130 (1.5%)
EQ (X-band) 122/130 (93.8%) 6/130 (4.6%) 2/130 (1.5%)
Table 6: Listening study - comparative evaluation results - 13 participants × 10 samples.
The comparative evaluation against existing baselines demonstrates SonicMaster’s superior performance across both reverb and EQ degradation categories (Table 6). For reverb artifacts, participants overwhelmingly preferred SonicMaster over Mel2Mel + Diffwave (Kandpal et al., 2022), selecting our method in 128 out of 130 total comparisons (10 samples × 13 participants), with Mel2Mel + Diffwave chosen only twice. In the EQ category, SonicMaster achieved similarly strong results with 122 out of 130 preferences, while Mel2Mel + Diffwave received 6 votes and Text2FX (Chu et al., 2025) garnered 2 votes. These results underscore SonicMaster’s effectiveness in addressing both spatial acoustic degradations and spectral imbalances compared to specialized baseline methods.
@@ A.2 Genre tags
A.3 Degradation functions
@@
A.4 Spectrogram examples
We visualize time–frequency structure in spectrograms to provide qualitative evidence of restoration behavior. Each figure shows the clean reference, the degraded input (e.g., reverberation-induced smearing or clipping distortion), and the output of SonicMaster. Figures 3, 4, 5, 6, 7, and 8 compare clean, degraded, and restored spectrograms across selected scenarios (reverb, clipping, microphone transfer function, and clarity EQ).

### v3

@@ 4.1 Baselines and training setup
SonicMaster (Ours) 0.095 0.754 0.380 7.627 0.121 1.251 0.368 7.477 0.101 1.002 0.374 7.552
Table 3: Objective Scores: FAD ($\downarrow$), KL ($\downarrow$), SSIM ($\uparrow$), and PQ ($\uparrow$). KL values are multiplied by 1000 for readability. Bold indicates best performance (excluding ground truth reference).
Evaluation is conducted along two orthogonal axes. (i) Global perceptual fidelity is quantified with FAD on CLAP embeddings (elizalde2023clap), Kullback–Leibler divergence (KL), structural similarity (SSIM) on 128-bin mel-spectrograms, and the Production Quality (PQ) score from the Audiobox Aesthetics toolbox (tjandra2025meta). (ii) Degradation-specific restoration efficacy is measured by average absolute error reduction: for every degraded clip in a $7000$ clip test set, we compute the relevant (based on the degradation deployed) artefact-aware metric against its clean counterpart from a $1000$ sample reference set, then recompute the metric after SonicMaster processing; the relative decrease indicates how closely each model variant approaches the ground-truth.
For X-band EQ and microphone-TF degradations, we compute the spectral balance over nine frequency bands and report their cosine distance. All other EQ effects are scored by the energy ratio between the affected band and the full spectrum. Compression is measured as the standard deviation of frame-level RMS (2048-sample frames, 1024 hop); punch as the mean onset-envelope value (librosa.onset.onset_strength). Because RT60 estimates are unreliable on dense mixes, reverb is assessed via the Euclidean distance of modulation spectra. Clipping uses spectral flatness; volume, the global RMS; and stereo width, the RMS ratio of the mid and side signals, $\mathrm{RMS}\!\left[\frac{L-R}{2}\right]/\mathrm{RMS}\!\left[\frac{L+R}{2}\right]$. We report the average absolute error value (GT vs inferred sample) of all the metrics except where mentioned differently (X-band, microphone-TF, and reverb). Details of each metric are described in Appendix A.5.
We presented listeners with $43$ audio sample pairs – degraded inputs and SonicMaster outputs – to rate, consisting of $2$ pairs for each degradation function ($2\times 19=38$ single degraded samples), $3$ pairs of double and $2$ pairs of triple degraded samples. Using a $7$-point Likert Scale, listeners were to rate: 1) The extent of improvement from the input to SonicMaster output represented by the text prompt (Text relevance), 2) audio quality of input (Quality1), 3) audio quality of the inferred SonicMaster sample (Quality2), 4) consistency and fluency of the inferred sample (Consistency), and 5) preference between the two samples, where 1 represents full preference of the ground truth degraded input, and 7 represents the SonicMaster inferred sample (Preference). The study was attended by 12 listeners (7 music experts and 5 Music Information Retrieval researchers).
Furthermore, to benchmark against existing methods, we conducted an additional study with 20 participants comparing SonicMaster against Text2FX (chu2025text2fx), Text2FX-directional, and Mel2Mel + Diffwave (kandpal2022music) on 20 randomly selected samples from our test set. The evaluation included 10 samples with X-band EQ degradation and 10 with reverberation artifacts. Note that Text2FX and Text2FX-directional are limited to EQ effects as their reverb effect is only additive, thus excluded. Since the baseline methods’ evaluation sets are not publicly available, we performed this comparison exclusively on our curated test data.
5 Results
5.1 Objective Evaluation
Degradation-Specific Performance: Tables 1 and 2 demonstrates SonicMaster’s superiority over baselines of Text2FX in EQ, and WPE/HPSS in Reverb. SonicMaster improves in all categories when compared to the degraded and reconstructed inputs. Furthermore, the reconstructed input metrics are overall slightly worse (with exceptions) than those of the ground truth degraded inputs.
@@ 5.1 Objective Evaluation
Perceptual Quality Assessment: Table 3 reveals SonicMaster outperforms the degraded inputs in both PQ and KL. FAD is marginally higher than that of the degraded audio, yet markedly lower than the reconstructed baseline. Furthermore, SonicMaster achieves a significant increase in PQ, almost reaching the level of ground truth mastered reference. In SSIM, SonicMaster exhibits lower scores than degraded inputs but achieves superior performance compared to the reconstruction baseline.
@@
Comparison with removal models: While models such as DPTNet chen2020dual, UMX stoter2019open, DCUNet choi2018phase, TCN rethage2018wavenet; steinmetz2021efficient, and HDemucs defossez2021hybrid focus on effect removal with minimal alteration rice2023general (best baseline: 20.08 dB for Dynamics, 13.59 dB for Reverb), SonicMaster performs text-guided mastering that applies intentional tonal and dynamic shaping. All baselines are trained following the RemFX protocol rice2023general using effect-specific supervision with L1 + multi-resolution STFT losses, and evaluated on the official test split containing clean vs. effected pairs for each degradation type. We test SonicMaster on the same test set, focusing on the two degradation: Dynamics and Reverb used (rice2023general). This broader objective enables SonicMaster to reconstruct a more coherent musical structure as shown in Fig. 3, achieving substantially higher SI-SDR scores of $47.11$ dB (Dynamics) and $45.76$ dB (Reverb).
Figure 3: Comparison of SI-SDR scores ($\uparrow$) for Dynamics and Reverb removal.
5.2 Ablation Studies
Text Prompt Dependency: Inference without text prompts maintains comparable FAD, SSIM, and PQ but shows degraded KL divergence (0.917 vs. 0.696). Critical drops occur in Clip restoration (2.812 vs. 1.506) and Stereo processing (0.1416 vs. 0.1058), with elevated EQ errors. To further assess the text controllability, we shuffled the prompts inside the test set and ran inference. Results (Tables 1, 2, 3 show worse performance than when no prompt was given (KL 2.014, Clip 2.874), but still show large improvement over the degraded input. This confirms text conditioning enables targeted restoration rather than generic improvements.
Architecture Scaling Analysis: We observe interesting scaling dynamics. SonicMasterSmall performs comparably with SonicMasterLarge in all metrics, but slightly worse in Reverb, Clip, and Stereo. SonicMasterMedium (4MM-DiT/12DiT) performs slightly better than the SonicMasterSmall, but still lacks behind SonicMasterLarge in Clip. SonicMasterMedium (6MM-DiT/6DiT) performs the worst out of all variants across all metrics. See Appendix A.7.
Audio Conditioning Duration: We evaluated SonicMasterLarge with different conditioning lengths (5s, 10s, 15s), finding comparable performance across configurations. The 10-second setting balances computational efficiency with temporal overlap for long-form processing. See Appendix A.7.
Conditioning Strategy Analysis: The no-conditioning variant achieves optimal Boom correction (0.0658 absolute error) but poor Clip restoration (2.055 vs. standard variants), highlighting multi-modal guidance importance for challenging tasks. More details in Appendix A.7.
Long-Form Audio Evaluation: Full-song evaluations confirm SonicMaster’s effectiveness, with substantial improvements in EQ-related metrics (Table 1) and most degradation functions (Table 2). Reverberation metrics show mixed results, likely due to increased complexity of spatial processing in extended musical contexts where room acoustics interact with diverse instrumental timbres and dynamic variations. SSIM and FAD decrease compared to degraded inputs, except for FAD in multi-degradation samples, indicating SonicMaster’s ability to handle compound degradations.
Figure 4: Listening study - SonicMaster’s performance on specific degradations – MOS 95% CI
@@ 5.3 Piano Recordings Evaluation
To test SonicMaster generalization, we evaluate historical solo piano pieces [^4] http://research.spa.aalto.fi/publications/papers/dafx-babe2/ using established baselines: LTAS-EQ, BEHM-GAN (9829821) model for bandwidth extension, and BABE/BABE-2 diffusion-based generative equalizers (moliner2024blind; moliner2023diffusion). BABE-2 represents a state-of-the-art specialized method for old recordings, it uses a diffusion prior to restore lost high frequencies and remove coloration, and has shown impressive improvements in archival music (moliner2024blind). Despite lacking domain-specific training, SonicMaster came surprisingly close to these specialized baselines (Table 4. In objective evaluations, SonicMaster restored samples achieved a PQ of $6.93$, nearly matching the $7.05$ obtained by BABE-2.
@@ 5.4 Subjective Evaluation
Figure 4 shows results of the first listening study. Text relevance ratings are highest in the Amplitude (6.21), Stereo (5.75), and Reverb (5.36) categories, indicating effective declipping, volume increase, expansion of the stereo image, and dereverberation. These three categories also show the highest consistency and preference ratings. The Dynamics and Amplitude categories show the biggest improvement in quality. EQ shows the fourth-best text relevance, but the worst preference ratings. This likely reflects the nature of some EQ effects being more stylistic or difficult to notice (e.g., airiness, boominess). Overall, SonicMaster samples are rated higher in quality compared to inputs and preferred across the board. A paired $t$-test on Quality1 and Quality2 ratings shows statistically significant differences ($p<0.05$ for Stereo, $p<0.01$ for the rest) in all categories except EQ.
The comparative evaluation against existing baselines demonstrates SonicMaster’s superior performance across both reverb and EQ degradation categories (Figure 5). For reverb artifacts, participants overwhelmingly preferred SonicMaster over Mel2Mel + Diffwave (kandpal2022music), selecting our method in 191 out of 200 total comparisons (10 samples $\times$ 20 participants), with Mel2Mel + Diffwave chosen only twice. In the EQ category, SonicMaster achieved similarly strong results with 180 out of 200 preferences, while Mel2Mel + Diffwave received 13 votes, Text2FX (chu2025text2fx) garnered 4 votes and Text2FX-directional generated 3. These results show SonicMaster’s effectiveness in addressing both spatial acoustic degradations and spectral imbalances.
Figure 5: Comparative Listening Study Results ($N=20$ participants $\times$ 10 samples per category).
@@ 7 Conclusion
Reproducibility Statement
We shall publicly release the implementation of model training, inference, evaluation, as well as dataset upon acceptance. We also mention the hyperparameters in the appendix.
@@ A.2 Genre tags
A.3 Text2FX-Directional Baseline for the EQ Task
For the equalization (EQ) experiments, we include the Text2FX-Directional method chu2025text2fx as a text-guided audio transformation baseline. Although Text2FX is not a restoration model, SonicMaster is instruction-conditioned; therefore, a text-conditioned FX optimizer offers a meaningful point of comparison for evaluating how well different systems follow natural-language EQ instructions.
A.3.1 Directional Loss Formulation
Text2FX-Directional uses CLAP audio/text embeddings to align the change in audio embedding with the semantic direction defined by a target prompt and a contrast prompt. Let $f_{a}$ and $f_{t}$ denote the CLAP audio and text encoders, and let $g(x;\theta)$ be a differentiable 6-band parametric EQ (dasp-pytorch). Given degraded audio $x_{\text{deg}}$ and prompts $t_{1}$ (contrast) and $t_{2}$ (target), we define:
$A_{1}$ $=f_{a}(x_{\text{deg}}),$
$A_{2}(\theta)$ $=f_{a}\bigl(g(x_{\text{deg}};\theta)\bigr),$
$T_{1}$ $=f_{t}(t_{1}),$
$T_{2}$ $=f_{t}(t_{2}).$
The method encourages the audio embedding to move from $A_{1}$ to $A_{2}$ in the same direction as the text embedding moves from $T_{1}$ to $T_{2}$. Let
$d_{a}(\theta)=\frac{A_{2}(\theta)-A_{1}}{\lVert A_{2}(\theta)-A_{1}\rVert_{2}},\qquad d_{t}=\frac{T_{2}-T_{1}}{\lVert T_{2}-T_{1}\rVert_{2}}.$
The directional loss is then:
$\mathcal{L}_{\text{dir}}(\theta)=1-\cos\bigl(d_{a}(\theta),d_{t}\bigr).$
We follow the optimization settings of chu2025text2fx: 600 Adam iterations (learning rate $1\times 10^{-2}$), standard-normal parameter initialization, and a random circular time shift at each step to avoid fixation on audio content.
A.3.2 Prompt and Contrast-Prompt Construction
Our EQ dataset contains natural-language instructions rather than the short adjectives used in chu2025text2fx. To maintain the $T_{1}\rightarrow T_{2}$ structure required by the directional loss, we construct a semantically opposite contrast prompt for each instruction using GPT with a constrained template (“write the opposite EQ action”) and manual verification.
Examples used in our EQ evaluation include:
  • Clarity / Treble Boost:
    • Target prompt ($T_{2}$): “Increase the clarity of this song by emphasizing treble frequencies.”
    • Contrast prompt ($T_{1}$): “Decrease the clarity of this song by softening or reducing the treble frequencies and making it sound more dull and muffled.”
  • Boominess / Low-End Enhancement:
    • Target prompt ($T_{2}$): “Add weight and depth to the bottom end.”
    • Contrast prompt ($T_{1}$): “Do the opposite of the following instruction: Add weight and depth to the bottom end.”
  • Mic / Narrow-Band Coloration:
    • Target prompt ($T_{2}$): “Balance the EQ, please.”
    • Contrast prompt ($T_{1}$): “Do the opposite of the following instruction: Balance the EQ, please.”
These pairs ensure that Text2FX-Directional receives properly opposed EQ semantics while matching the full-sentence instruction style of our enhancement dataset.
A.3.3 Purpose of This Baseline
Text2FX-Directional does not use the clean reference audio during optimization; thus it is not evaluated as a restoration model. Instead, we include it as a text-conditioned equalization baseline that evaluates: How well can a CLAP-guided, single-instance EQ optimizer follow the same natural-language instructions given to SonicMaster? This provides a fair, instruction-aligned comparison for EQ-specific transformations under identical textual guidance.
A.4 Degradation functions
@@
Peak normalisation of tracks: In case of no intentional clipping, “hidden clipping”, or a low volume degradation being used, all degraded versions of the SonicMaster dataset are normalised to a peak amplitude $y_{peak}$ drawn from a uniform distribution $y_{peak}\sim U(0.8,1.0)$, track is then normalised as:
$x_{norm}=\frac{x}{max(abs(x))}\times y_{peak}$
@@
A.5 Evaluation metrics details
To evaluate SonicMaster’s ability to deal with each of the 19 proposed degradations, we use a set of evaluation metrics as follows in this section. For all the metrics, except for X-band EQ, microphone transfer function, and all reverb options, we report absolute errors, i.e., the absolute value of difference of ground truth (GT) and inferred sample metric values:
$AbsError_{metric}=\bigl|metric_{ground\_truth}-metric_{inferred}\bigl|.$
EQ: The effect of all the EQ options, except for ”xband” and ”mic” is evaluated through absolute error of spectral energy ratio of two signals – the ground truth reference and the inferred signals. Spectral energy ratio (Spectral_ERSpectral\_ER) is computed as:
$Spectral\_ER=\frac{E_{band}}{E_{total}},$
where $E_{total}$ is the total energy of the signal, and $E_{band}$ is the signal’s energy in a spectral band given by the following boundaries $B$:
[…]
where $f_{s}$ stands for sampling rate.
The remaining two EQ functions of “xband” and “mic” are evaluated through a cosine distance of spectral balance of the ground truth reference and inferred signal. Spectral balance ($SB$) is calculated as a normalised energy profile in 9 pre-defined frequency bands:
$SB=\frac{[E_{1},E_{2},E_{3},E_{4},E_{5},E_{6},E_{7},E_{8},E_{9}]}{sum[E_{1},E_{2},E_{3},E_{4},E_{5},E_{6},E_{7},E_{8},E_{9}]}.$
The bands are given as:
[…]
The reported cosine distance is then gained as:
$cosine\_distance=1-cos(SB_{ground\_truth},SB_{inferred}).$
Amplitude: Clipping correction is evaluated through spectral flatness using the librosa.feature.spectral_flatness library function, which takes in a power spectrogram gained through STFT with n_fft=2048 and hop length = 512. The final metric for clipping is the absolute error of spectral flatness (GT vs inferred sample).
Volume is evaluated as the absolute error of the Root-Mean-Square (RMS) value.
Dynamics: Compression is evaluated as the standard deviation of the dynamic range (STD_DRSTD\_DR), given as:
$STD\_DR=std(RMS(\mathcal{F}_{H,L})),$
where $\mathcal{F}_{H,L}$ represents a set of waveform frames with length 2048 and hop length 1024 each. The final metric is the absolute error from the GT.
The “punch” is measured through transient strength by taking the mean value of the transient envelope gained from the librosa.onest.onset
strength
library function with default parameters.
Reverb: We evaluate the effect of dereverberation using modulation spectrum distance ($MSD$).
First, we get a set of temporal envelopes $E_{x}$ from input signal $x$:
$E_{x}^{(k)}(m)=\bigl|\mathrm{STFT}\{x\}(k,m)\bigr|,$
where $k$ indexes frequency bins and $m$ indexes time frames. Modulation spectrum $S_{x}^{(k)}(b)$ is then calculated using demeaned temporal envelopes:
$S_{x}^{(k)}(b)=\left|\mathrm{FFT}_{m}\!\left(E_{x}^{(k)}(m)-\frac{1}{M}\sum_{m^{\prime}=0}^{M-1}E_{x}^{(k)}(m^{\prime})\right)\right|_{b},\qquad b=0,\dots,B-1.$
where $b$ represents modulation bins.
Modulation spectra from all frequency bands are then stacked into a single vector:
$\mathbf{s}_{x}=\mathrm{vec}\!\left(S_{x}^{(k)}(b)\right),$
and $\ell_{2}$ normalized:
$\hat{\mathbf{s}}_{x}=\frac{\mathbf{s}_{x}}{\|\mathbf{s}_{x}\|_{2}+\varepsilon}.$
The MSD between two signals, in our case the GT reference $x_{GT}$ and relevant inferred sample $x_{infer}$, is given as Euclidean distance:
$MSD(x_{GT},x_{infer})=\left\|\hat{\mathbf{s}}_{x_{GT}}-\hat{\mathbf{s}}_{x_{infer}}\right\|_{2}.$
In code, this is realized with following parameters as:
import numpy as np
from scipy.spatial.distance import euclidean
from scipy.signal import stft
def modulation_spectrum_distance(x1, x2, fs=44100,
    n_fft=1024, hop_length=512, n_mod_bins=20):
    def get_modulation_spectrum(x):
        f, t, Zxx = stft(x, fs=fs, nperseg=n_fft, noverlap=n_fft - hop_length)
        mag = np.abs(Zxx)
        mod_spec = []
        for band in mag:
            envelope = band - np.mean(band)
            spectrum = np.abs(np.fft.fft(envelope))[:n_mod_bins]
            mod_spec.append(spectrum)
        mod_spec = np.array(mod_spec)
        mod_spec /= np.linalg.norm(mod_spec) + 1e-10
        return mod_spec.flatten()
    mod1 = get_modulation_spectrum(x1)
    mod2 = get_modulation_spectrum(x2)
    return euclidean(mod1, mod2)
Stereo: We measure the level of stereoness using stereo energy ratio (Stereo_ERStereo\_ER), computed as:
$Stereo\_ER=\frac{RMS(\frac{L-R}{2})}{RMS(\frac{L+R}{2})+10^{-10}}$ (3)
We report the absolute error of this metric.
A.6 Spectrogram examples
We visualize time–frequency structure in spectrograms to provide qualitative evidence of restoration behavior. Each figure shows the clean reference, the degraded input (e.g., reverberation-induced smearing or clipping distortion), and the output of SonicMaster. Figures 6, 7, 8, 9, 10, and 11 compare clean, degraded, and restored spectrograms across selected scenarios (reverb, clipping, microphone transfer function, and clarity EQ).